# Simplifying Traffic Anomaly Detection with Video Foundation Models

### v1

@@ 3.1 Task definition
In the general case, $\tau$ can be 1 and $f_{\theta}$ may also maintain an internal hidden state or operate in an autoregressive manner, explicitly conditioning on previously predicted outputs.
@@ 3.2 Evaluation Procedure
Prior work typically reports the Area Under the Receiver Operating Characteristic Curve (AUCROC) as the primary evaluation metric for TAD [48, 51, 24, 32, 25], and we adopt this metric when comparing to previous methods. However, handling data imbalance is especially important in TAD, so in our evaluations we use the Matthews Correlation Coefficient (MCC), which has been used in related work [42, 18, 37]. MCC takes into account all entries of the confusion matrix, including true negatives, and better reflects overall performance under class imbalance [7]. MCC at a given threshold is defined as:
@@
where TP, TN, FP, and FN denote true positives, true negatives, false positives, and false negatives, respectively. Note that MCC ranges from $-1$ (inverse prediction) to $1$ (perfect prediction), with 00 indicating random performance, but we show it in the range -100, 100 to improve readability.
@@
Beyond metric design, we implement a broader protocol focused on generalization and efficiency. We evaluate in-domain performance, out-of-domain performance, and computational cost.
@@ 3.3 Pre-trained Encoders
We select a range of recent ViFMs that represent various pre-training strategies, and apply them to the TAD task; see Tab. 1 for an overview of their pre-training strategies. When possible, we select variants pre-trained on Kinetics-400 [19] for consistency.
We include ViViT [1] as a baseline to represent fully-supervised pre-training. VideoMAE [38], MVD [45], and VideoMAE2 [44] are selected to evaluate progressively stronger variants of self-supervised pre-training from videos. We also assess InternVideo2 [46], which combines self-supervised learning from videos and weakly-supervised learning from multiple modalities, and is one of the leading models across numerous video benchmarks. Together, these ViT-based models cover a diverse range of pre-training strategies.
For completeness, we also include the fully-convolutional R(2+1)D model, pretrained in a fully-supervised manner, motivated by recent studies showing its competitive performance in the related task of collision anticipation [10].
@@ 3.4 Domain-Adaptive Pre-Training
To better align the Video ViT encoder with the driving domain, we adopt the Domain-Adaptive Pre-Training (DAPT) strategy, a simple method originally proposed in the field of natural language processing [16]. DAPT introduces an additional pre-training stage between generic pre-training and downstream fine-tuning, using unlabeled data from the target domain.
@@ 4 Experiments
DAPT-VideoMAE-L (ours) 88.4 84.2 88.5 86.6 304 M 1.80 GB 34
@@ 4.1 Experimental setup
Datasets.  We evaluate on DoTA [48] and DADA-2000 [13], two large-scale real-world driving anomaly datasets with temporal and frame-level annotations. DAPT uses Kinetics-700 [5], BDD100K [49], and CAP-DATA [14], described in detail in the Supplementary.
Model input.  All Video ViTs and R(2+1)D are trained on sliding windows of size 224×224×16 at 10 FPS (1.5s temporal context) by default. For InternVideo2, which uses tubelets of size 1, we use 224×224×8 at 5 FPS to match the same duration. MOVAD processes videos frame-by-frame at resolution 640×480.
Fine-tuning.  With all Video ViTs and R(2+1)D, we closely follow the VideoMAE fine-tuning recipe for HMDB51. We train for 50 epochs (5 warmup), with 50K randomly sampled examples per epoch and a batch size of 56. For VidNeXt variants, and MOVAD, we follow the original training settings.
Domain-adaptive pre-training (DAPT).  We apply the VideoMAE pre-training strategy [38], masking 75% of tokens, using MSE loss on masked tokens only. Training uses a batch size of 800 and 1M samples per epoch, with 12 epochs. We explore DAPT on three domains: (a) Kinetics-700, (b) BDD100K (normal driving), and (c) BDD100K + CAP-DATA (abnormal driving), with dataset mixing ratios detailed in the Supplementary.
@@ 4.2 Can an encoder-only model outperform specialized TAD methods?
To answer this question, we evaluate models along three critical axes: classification performance, generalization, and efficiency, as shown in Tab. 2. We select a range of recent top-performing methods proposed for the TAD task. Among encoder-only models, we apply the R(2+1)D [39] model and different sizes of VideoMAE pre-trained Video ViTs (with DAPT, see Sec. 4.4). The results show that these Video ViTs consistently strike a strong balance, demonstrating a good combination of predictive accuracy, generalization across domains, and computational efficiency. Notably, strongly pre-trained Video ViTs achieve the highest AUCROC scores across both DoTA and DADA-2000 datasets, both in-domain and in cross-dataset evaluation, while being highly efficient with a low memory footprint. In contrast, specialized TAD-specific models not only demonstrate lower classification performance but also incur substantially higher computational costs and latency. R(2+1)D and ResNet+NST, while being highly efficient, fall short in predictive quality. This confirms that we can outperform specialized, multi-component TAD methods with a simple encoder-only model by applying a Video ViT with strong pre-training.
@@ 4.3 What pre-training is better for TAD?
We investigate how general video recognition performance and pre-training strategies relate to downstream performance on TAD. We evaluate a range of publicly available Video ViT models of several sizes, using their Top-1 accuracy on Kinetics-400 [19] and Something-SomethingV2 [15] alongside AUCMCC on DoTA[48] and DADA-2000 [13].
Results are summarized in Tab. 3, from which we observe two key trends. First, we find that for TAD the MAE pre-training objective dominates: MAE-pre-trained models (VideoMAE) and their distilled variants (VideoMAE2, MVD) achieve the highest AUCMCC within each size tier, even when they do not have the highest classification accuracy on general benchmarks. Second, classification accuracy on general benchmarks is not representative of TAD performance: ViViT-B, despite matching VideoMAE on Kinetics-400, demonstrates significantly lower AUCMCC, and InternVideo2, state-of-the-art on general benchmarks, also underperforms on TAD. These findings suggest that the representations which are beneficial for general video classification may not align well with those needed for TAD. In particular, TAD appears to benefit more from dense representations that emphasize fine-grained temporal irregularities rather than the coarse semantic categories typically targeted by general video recognition models.
@@
Model Variant Type K400 SthSthV2 DoTA D2K DoTA D2K
@@
Table 3: Comparing Video ViT pre-trainings. In contrast to general video classification benchmarks (K400, SthSthV2), fully- and weakly-supervised pre-training are less effective for TAD benchmarks. Self-supervised pre-training performs best for TAD (DoTA, D2K). FSL: fully-supervised; WSL: weakly-supervised; SSL: self-supervised learning; K400: Kinetics-400 [19]; SthSthV2: Something-SomethingV2 [15]; D2K: DADA-2000 [13].
@@ 4.4 Domain-Adaptive Pre-Training (DAPT)
Larger ViFMs can be pre-trained on a larger scale and, as a result, exhibit better out-of-the-box generalization across domains, while smaller models have shown to benefit less from longer pre-training due to their limited capacity and faster saturation [50]. Therefore, we expect that domain adaptation can help better utilize the capacity of the smaller but at the same time more efficient and faster models. Given that MAE pre-training proves especially effective for TAD, and unlabeled driving data is available in abundance, we investigate whether small-scale self-supervised DAPT with MAE can be an effective and efficient way to scale the performance of smaller models. We initialize a ViT model with VideoMAE pre-trained weights and perform several epochs of additional pre-training with the VideoMAE objective on in-domain data. Compared to the original $\sim$192K training steps with batch size 2048, we use only 15K steps with batch size 800.
@@
Figure 3: DAPT scaling across different model sizes. Smaller models benefit more. S: small, B: base, L: large variant of the Video ViT.
To disentangle the impact of domain relevance from that of additional pre-training, we conduct an ablation study on the data used for DAPT. Specifically, we adapt a model, pre-trained on a general human activity dataset, using three types of unlabeled video data: (1) the original, general pre-training domain used by VideoMAE, which is not related to the TAD task, (2) normal ego-centric driving, and (3) ego-centric driving mixed with anomalies. This setup allows us to evaluate whether the observed improvements stem from domain alignment or from simply continuing generic pre-training. As shown in Tab. 4, adaptation with domain-relevant data (both normal and abnormal driving) consistently improves generalization and data efficiency, while additional pre-training on the original, generic domain yields no notable gains. Interestingly, pre-training on normal driving videos is sufficient, and mixing in data with driving anomalies does not provide further improvements. We conclude that small-scale self-supervised DAPT is a simple and effective way to improve the performance and generalization of smaller Video ViTs for TAD, which does not necessarily requires rare anomaly data.
DoTA AUCMCC, D2K AUCMCC
@@
Driving DAPT 54.8 +2.7 46.8 +3.0 52.0 +3.9 49.7 +3.1
+ anomalies 54.9 +2.8 46.9 +3.1 51.9 +3.8 49.7 +3.1
@@
Figure 4: Qualitative examples for the effect of DAPT. Predicted anomaly‐scores of VideoMAE (top: S, bottom: B) w/ and w/o DAPT.
@@ 5 Discussion
In this work, we show that with stronger pre-training, an encoder-only Video Vision Transformer outperforms all prior Traffic Anomaly Detection models while also being significantly more efficient. However, it remains an open question whether the additional components introduced in earlier work become redundant as pre-training scales, as shown in related perception tasks [21], or whether they still provide complementary benefits.
We use Video Masked Autoencoders [38, 44] as a simple and effective form of masked video modeling (MVM). Approaches that predict in latent space, such as V-JEPA [3, 2], may offer further gains by avoiding the limitations of noisy pixel-level reconstruction.
@@ 6 Conclusion
Ego-centric Traffic Anomaly Detection (TAD) is a challenging task that requires modeling motion dynamics and agent interactions. While most recent methods for TAD rely on complex, multi-component architectures, we show that a simple encoder-only design using a plain Video Vision Transformer (ViT) with strong self-supervised pre-training is not only more efficient, but also more effective and generalizable. Building on this, Domain-Adaptive Pre-Training (DAPT) offers a label-free and data-efficient way to further boost performance, particularly for smaller models. These findings highlight the strength of learned inductive biases from large-scale pre-training as an alternative to manually crafted architectural complexity, a principle to which TAD is no exception.
Our experiments further demonstrate that Masked Video Modeling (MVM) is the most effective pre-training strategy for TAD, in contrast to standard video classification tasks. This suggests that different video tasks may benefit from pre-training objectives tailored to their downstream requirements. While TAD is a crucial task in autonomous driving (AD), other AD tasks may align more closely with conventional action recognition. This motivates further research into a universally effective video pre-training strategy, evaluated by its generalization across diverse AD tasks. We hope our findings provide a foundation for future work in this direction.

### v2

@@ 3.1 Task definition
In the general case, $\tau$ can be 1 and $f_{\theta}$ may use recurrence or autoregressive conditioning on past predictions.
@@ 3.2 Evaluation Procedure
Prior work typically reports the Area Under the Receiver Operating Characteristic Curve (AUCROC) as the primary evaluation metric for TAD [54, 57, 24, 35, 25], and we adopt it as well for comparison with previous methods. However, because handling data imbalance is especially important in TAD, we follow related work [48, 16, 42] and also report the Matthews Correlation Coefficient (MCC). MCC considers all entries of the confusion matrix, including true negatives, and better reflects overall performance under class imbalance [5]. MCC at a given threshold is defined as:
@@
where TP, TN, FP, and FN denote true positives, true negatives, false positives, and false negatives, respectively. Note that MCC ranges from 1-1 (inverse prediction) to $1$ (perfect prediction), with 0 indicating random performance, but we show it in the range -100, 100 to improve readability.
@@
Beyond metric design, we adopt a broader evaluation protocol that emphasizes both generalization and efficiency. We assess in-domain and out-of-domain performance, as well as efficiency based on parameter count, peak GPU memory usage, and frames per second (FPS).
@@ 3.3 Pre-trained Encoders
We select a range of recent Video ViT-based ViFMs that represent various pre-training strategies, and apply them to the TAD task; see Tab. 1 for an overview of their pre-training strategies. When possible, we select variants pre-trained on Kinetics-400 [17] for consistency. For completeness, we also include R(2+1)D [45], a fully-convolutional encoder, motivated by recent studies showing its competitive performance in the related task of collision anticipation [8].
We include ViViT [1] and fully-convolutional R(2+1)D [45] to represent fully-supervised pre-training. VideoMAE [44], MVD [51], and VideoMAE2 [50] are selected to evaluate progressively stronger variants of self-supervised pre-training from videos. Among weakly-supervised methods, we assess UMT [22] and SMILE [43], both of which align video tokens with CLIP [36] supervision. Finally, InternVideo2 [52] combines self-supervised learning from videos and weakly-supervised learning from multiple modalities, and is one of the leading models across numerous video benchmarks. Together, these models cover a diverse range of pre-training strategies.
@@ 3.4 Domain-Adaptive Pre-Training
To better align the Video ViT encoder with the driving domain, we adopt the Domain-Adaptive Pre-Training (DAPT) strategy, a simple method originally proposed in the field of natural language processing [13]. DAPT introduces an additional pre-training stage of smaller scale between generic pre-training and downstream fine-tuning, using unlabeled data from the target domain.
@@ 4 Experiments
DAPT-VideoMAE-L (ours) 88.4 85.0 88.5 86.6 304 M 1.80 GB 34
@@ 4.1 Experimental setup
Datasets.  We evaluate on DoTA [54] and DADA-2000 [10], two large-scale real-world driving anomaly datasets with temporal and frame-level annotations. We also manually refine 1% of annotations of the DoTA dataset and refer to this variant as DoTAref, more details in Appendix D. DAPT uses Kinetics-700 [3], BDD100K [55], and CAP-DATA [11].
Model input.  All Video ViTs and R(2+1)D are trained on sliding windows of size 224×224×16 at 10 FPS (1.5s temporal context) by default. For InternVideo2 and UMT, which use tubelets of size 1, we use 224×224×8 at 5 FPS to match the same duration. MOVAD processes videos frame-by-frame at resolution 640×480.
Fine-tuning.  With all Video ViTs, R(2+1)D, and VidNeXt variants, we closely follow the VideoMAE fine-tuning recipe for HMDB51. We train for 50 epochs (5 warmup), with 50K randomly sampled examples per epoch and a batch size of 56. For MOVAD, we follow the original training settings.
Domain-adaptive pre-training (DAPT).  We apply the VideoMAE pre-training strategy [44], masking 75% of tokens, using MSE loss on masked tokens only. Training uses a batch size of 800 and 1M samples per epoch, with 12 epochs. We explore DAPT on three domains: (a) Kinetics-700, (b) BDD100K (normal driving), and (c) BDD100K + CAP-DATA (a mix of normal and abnormal driving). More details can be found in Appendix A.
@@ 4.2 Can an encoder-only model outperform specialized TAD methods?
To answer this question, we evaluate models along three critical axes: classification performance, generalization, and efficiency, as shown in Tab. 2. We select a range of recent top-performing methods proposed for the TAD task. Among encoder-only models, we apply the R(2+1)D [45] model and different sizes of VideoMAE pre-trained Video ViTs (with DAPT, see Sec. 4.4).
The results show that these Video ViTs consistently strike a strong balance, demonstrating a good combination of predictive accuracy, generalization across datasets, and computational efficiency. Notably, Video ViTs with advanced pre-training achieve the highest AUCROC scores across both DoTA and DADA-2000 datasets, both in-domain and in cross-dataset evaluation, while being highly efficient with a low memory footprint. In contrast, specialized TAD-specific models not only demonstrate lower classification performance but also incur substantially higher computational costs and latency. R(2+1)D and ResNet+NST, while being highly efficient, fall short in predictive quality. This confirms that we can outperform specialized, multi-component TAD methods with a simple encoder-only model by applying a Video ViT with advanced pre-training. A more detailed comparison is provided in Appendix C.
@@ 4.3 What pre-training is better for TAD?
We evaluate a range of publicly available Video ViT models of several sizes, using their Top-1 accuracy on the general benchmarks Kinetics-400 [17] and Something-SomethingV2 [12] alongside AUCMCC and MCC@0.5 on the TAD benchmarks DoTA [54] and DADA-2000 [10]. Results are shown in Tab. 3, from which we observe three key trends.
First, we find that for TAD, self-supervised pre-training, i.e. the Masked Video Modeling (MVM) objective dominates: models, pre-trained with the Masked Autoencoding (MAE) approach (VideoMAE), and their distilled variants (VideoMAE2, MVD) achieve the highest AUCMCC within each size tier. Additional experiments on different MVM objectives can be found in Appendix B. Second, models that employ weakly-supervised pre-training (UMT, SMILE and InternVideo2) perform worse on TAD, indicating that vision-language supervision may not transfer well to fine-grained temporal anomaly understanding. Finally, classification accuracy on general benchmarks is not representative of TAD performance: ViViT-Base, despite matching VideoMAE on Kinetics-400, demonstrates significantly lower AUCMCC, and InternVideo2, state-of-the-art on general benchmarks, clearly underperforms on TAD. This suggests that the representations that are beneficial for general video classification may not align well with those needed for TAD. In particular, TAD appears to benefit more from dense representations that emphasize fine-grained temporal irregularities rather than the coarse semantic categories typically targeted by general video recognition models.
@@
Model Variant Type K400 SthSthV2 DoTAref D2K DoTAref D2K
@@
Table 3: Comparing Video ViT pre-training strategies. In contrast to general video classification benchmarks (K400, SthSthV2), fully and weakly supervised pre-training are less effective for TAD (DoTAref, D2K), while self-supervised pre-training yields the best performance. FSL: fully-supervised; WSL: weakly-supervised; SSL: self-supervised learning; K400: Kinetics-400 [17]; SthSthV2: Something-SomethingV2 [12]; D2K: DADA-2000 [10].
Figure 3: DAPT scaling across different model sizes. Smaller models benefit more. S: Small, B: Base, L: Large variants of the Video ViT.
@@ 4.4 Domain-Adaptive Pre-Training (DAPT)
Larger ViFMs can be pre-trained on a larger scale and, as a result, exhibit better out-of-the-box generalization across domains, while smaller models have shown to benefit less from longer pre-training due to their limited capacity and faster saturation [56]. Therefore, we expect that domain adaptation can help better utilize the capacity of smaller, yet more efficient and faster models. Given that MAE pre-training proves especially effective for TAD, and unlabeled driving data is available in abundance, we investigate whether small-scale self-supervised DAPT with MAE can be an effective and efficient way to scale the performance of smaller models. We initialize a ViT model with VideoMAE pre-trained weights and perform several epochs of additional pre-training with the VideoMAE objective on in-domain data. Compared to the original $\sim$192K training steps with batch size 2048, we use only 15K steps with batch size 800.
@@
DoTAref, AUCMCC D2K, AUCMCC
@@
Driving DAPT 54.9 +2.8 47.0 +3.2 52.1 +4.0 49.7 +3.1
+ anomalies 55.0 +2.9 47.1 +3.3 52.0 +3.9 49.8 +3.2
@@
Figure 4: Qualitative examples for the effect of DAPT. Predicted anomaly‐scores of VideoMAE with and without DAPT (top: Video ViT-Small, bottom: Video ViT-Base).
To disentangle the impact of domain relevance from that of additional pre-training, we conduct an ablation study on the data used for DAPT. Specifically, we adapt a VideoMAE model, pre-trained on a general human activity dataset, using three types of unlabeled video data: (1) the original, general pre-training domain, which is not related to the TAD task, (2) normal ego-centric driving, and (3) ego-centric driving mixed with anomalies. This setup allows us to evaluate whether the observed improvements stem from domain alignment or from simply continuing pre-training. As shown in Tab. 4, adaptation with domain-relevant data (both normal and abnormal driving) consistently improves in-domain performance and generalization, while additional pre-training on the original, generic domain yields no notable gains. Interestingly, pre-training on normal driving videos is sufficient, and mixing in data with driving anomalies does not provide any further improvement. We conclude that small-scale self-supervised DAPT is a simple and effective way to improve the performance and generalization of smaller Video ViTs for TAD, without the need to collect rare anomalous examples.
@@ 5 Discussion
In this work, we show that with advanced pre-training, an encoder-only Video Vision Transformer outperforms all prior Traffic Anomaly Detection models while also being significantly more efficient. However, it remains an open question whether the additional components introduced in earlier work become redundant as pre-training scales, as shown in related perception tasks [19], or whether they still provide complementary benefits.
@@ 6 Conclusion
Ego-centric Traffic Anomaly Detection (TAD) is a challenging task that requires modeling motion dynamics and agent interactions. While most recent methods for TAD rely on complex, multi-component architectures, we show that a simple encoder-only design using a plain Video Vision Transformer with strong self-supervised pre-training is not only more efficient, but also more effective and generalizable. Building on this, Domain-Adaptive Pre-Training offers a label-free and compute-efficient way to further boost performance, particularly for smaller models. These findings highlight the strength of learned inductive biases from large-scale pre-training as an alternative to manually crafted architectural complexity, a principle to which TAD is no exception.
Our experiments demonstrate that Masked Video Modeling is the most effective pre-training strategy for TAD, in contrast to general video classification tasks. This suggests that different video tasks may benefit from different pre-training objectives tailored to their downstream requirements. While TAD is a crucial task in autonomous driving (AD), other AD tasks may align more closely with conventional action recognition. This motivates further research into a universally effective video pre-training strategy, evaluated by its generalization across diverse AD tasks. We hope our findings provide a foundation for future work in this direction.
@@ Acknowledgements
This work was funded by the Horizon Europe programme of the European Union, under grant agreement 101076754 (project AITHENA).